OpenStudy (anonymous):

help! medal! what are the zeros of the function? f(x)=x^3+6x^2+8x

OpenStudy (anonymous):

first take x common

OpenStudy (anonymous):

x(x^2+6x+8)=x(x^2+4x+2x+8) =x(x(x+4)+2(x+4)) =x(x+4)(x+2)

OpenStudy (anonymous):

so it would be 0,2,4?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

0 -2 -4

OpenStudy (anonymous):

oh ok thanks!